On 30 September 2019, office desk purchased on 30 April 2017 at a cost price of R5 200 was sold for R4 800 cash. A replacement equipment was purchased at a cost of R43 000 on 1 November 2019. The total depreciation (correctly calculated) for the year on office equipment amounted to R15 200. 2. Land and buildings were acquired on 1 April 2019. 3. On 28 February 2019, the business purchased a new tipper truck at a cost of R101 000. 4. It is the accounting policy of Tera Trader to provide depreciation as follows: Vehicles: According to the straight-line method, at 20% per annum, Office equipment: According to the diminishing-balance method, at 18% per annum, Buildings: According to the straight-line method, at 2% per annum. QUESTION 1 Which one of the following alternatives represents the correct amount of accumulated depreciation on office equipment which was sold on 30 September 2019? 1. R507 2. R1 448 3. R874 4. R1 955
1
Expert's answer
2020-02-26T09:52:46-0500


In connection with the acquisition of an office desk on April 30, 2017 for 5200, this amount of 5200 will be the amount from which further depreciation (initial cost) will be calculated. Since depreciation is calculated using the decreasing balance method, each time the calculation is made of decreasing residual value due to depreciation of equipment:


  1. 5200 * 18%/100%= 936 - depreciation from May 1, 2017 to May 1, 2018
  2. 5200-936 = 4264 - residual value after the first year of use
  3. 4264 * 18%/100%= 767.52 - depreciation from May 1, 2018 to May 1, 2019
  4. 4264-767.52 = 3496.48 - residual value after the second year of use
  5. 3496.48 * 18%/100%= 629.36 - depreciation for 2019, 629.36 * 1,5%(18% / 12) = 9.44 - a decrease in monthly depreciation in each month in 2019
  6. 629.36 / 12 * 5 = 262.24 - depreciation from May 1 to September 30, 2019
  7. 936 + 767.52 + 262.24 = 1965.76 - depreciation for the entire period of use without monthly reduction for 2019
  8. 1965.76-9.44 = 1955 - depreciation for the entire period of use


answer 4